Solve by completing the square.
x^2 = 5x + 2

Answer by Nate(3500)

x^2 = 5x + 2
x^2 - 5x = 2
(x - 2.5)^2 = 2 + 2.5^2 = 8.25
x - 2.5 = +- sqrt(8.25)
x = 2.5 +- sqrt(8.25)